# references.bib

### v2

@@ 3.2.2 Tomato growth conditions in hydroponic culture
Within these buckets, the nutrient mixture was aerated and mixed by introducing air supplied by a blower (YASUNAGA, AP-40P). The air was divided into eight segments, each channeled through silicone tubing and dispersed into the nutrient solution. The use of air stones ensures proper mixing and oxygenation. The buckets were equipped with lids containing center holes for the tomato stems to prevent algae growth in the nutrient solution.
@@ 3.2.3 Fertilizer concentration settings
The nutrient solution was prepared via hydroponic fertilizer (OAT Agrio, Fertilizer Series SA), and maintained at three different total dissolved solids (TDS) concentrations to induce various levels of fertilizer stress:
  • High concentration: 1500 ppm
  • Intermediate concentration: 1000 ppm
  • Low concentration: 500 ppm
This study employed an interventional research design to elucidate the causal relationship between fertilizer stress and tomato phenotypes. We set three fertilizer levels to investigate the presence or absence of intervention effects on tomato phenotypes and their dose-response relationships.
@@
Based on these findings, we set the low-concentration treatment at 500 ppm TDS. The intermediate- and high-concentration treatments were set at twice this level (1000 ppm) and three times this level (1500 ppm), respectively. Furthermore, to confirm that this range of concentrations did not exert excessively adverse effects on tomato growth, we conducted preliminary cultivation experiments. The results revealed no visually identifiable growth defects or plant death at 500 ppm. Additionally, even at 1500 ppm, no apparent growth inhibition due to increased osmotic pressure was observed. These results confirmed the suitability of the selected fertilizer concentration range for the research objectives.
@@ 3.2.4 Experimental design and management
The fertilizer plots were arranged in a randomized block design, as shown in Figure 7. During the preliminary experiment, the tomato plants were initially cultivated in the vacant areas depicted in Figure 7. To evaluate the effects of fertilizer level on outcome variables, such as the concentration of nitrate ions in tomato leaves, the density of trichomes, and yield, this randomized block design was employed to minimize the potential impact of spatial variations in environmental factors [fisher1966design].
Water was added to the buckets as needed when the nutrient solution water level decreased. Fertilizer was applied every two weeks, and nutrient solution concentrations were adjusted to maintain the desired TDS levels, which were measured via a TDS meter (EUTECH, PCSTestr35). The fertilizer used in the experiments closely matched the nutrient uptake ratios of the tomatoes, enabling the use of a closed hydroponic system without frequent nutrient solution replacement [ishihara2007effect]. However, during the cultivation period, the system was not maintained as a completely closed system; nutrient solutions were replaced whenever they became cloudy because algae or root debris was present.
During the cultivation period, two plants were lost before the compound leaves were harvested, leaving 20 plants available for inclusion in the study. Furthermore, an additional two plants were lost by the end of the fruit harvesting period, resulting in 18 plants being available for harvest yield assessment.
@@ 3.2.5 Compound leaf sampling and trichome transfer
The experiment involved collecting compound leaves from tomato plants. A total of 40 compound leaf samples were collected during two separate harvesting sessions in March and April 2023. The sampling process followed a systematic approach to ensure consistency and minimize potential biases. The samples were processed sequentially in ascending order of the number of tomato plant samples. These sample numbers were randomly assigned to different fertilizer treatment levels as part of the randomized block design, ensuring that the sampling order was not biased by treatment level.
The compound leaves were sampled 15 cm below the apical meristem of each plant. If no suitable compound leaf was available at this position, the plant was temporarily skipped, allowing for further growth. The sampling process continued iteratively until all the plants had been sampled. For trichome density measurements, we selected all leaflets larger than the opening of the measurement paper from each compound leaf.
The number of measurement papers used corresponded to the number of selected leaflets per compound leaf: This approach allows for an accurate assessment of trichome density distribution across each leaflet. Prior to image capture, glandular trichomes were transferred from the leaflet surface to the measurement paper. Each leaflet was carefully affixed to the adhesive side of the cellophane tape (Figure 8) and gently pressed with a finger to ensure complete contact between the leaf surface and the tape. After the leaflet was removed, the cellophane tape with the transferred trichomes was photographed alongside the measurement paper via a smartphone camera.
@@ 3.2.6 Image acquisition and dataset preparation
The measurement paper was photographed via the native camera app on a smartphone (OPPO, Reno A). This device has a 16-megapixel camera featuring an F1.7 aperture and a CMOS sensor (Sony, IMX398). The camera supports autofocus functionality, allowing for sharp and clear images even under varying conditions. For our experiments, the camera mode was set to automatically adjust the exposure, ISO sensitivity, and focus, leveraging the device’s built-in optimization algorithms.
To ensure a comprehensive dataset that accounts for various usage scenarios, we manually varied the distance and angle between the camera and the measurement paper. The distance from the camera sensor to the measurement paper ranged from 79 to 218 mm, simulating different user holding positions. The horizontal and vertical offset angles varied from -12.52 to 10.62 degrees and -9.04 to 8.13 degrees, respectively. This range of angles mimics slight misalignments that might occur during practical use of the kit. Furthermore, the measurement paper was positioned a sufficient distance from the background to ensure that the background was blurred and not clearly identifiable in the captured images.
The room used for photography had a north-facing window, which eliminated direct sunlight and provided diffuse natural light. The photographic sessions were conducted at different times of the day, with solar elevation angles ranging from -7.95 to 33.42 degrees. The smartphone’s built-in white LED was activated for auxiliary lighting, and the measurement paper was additionally illuminated by oblique light from the ceiling’s fluorescent light (as shown in Figure 4).
The image composition was carefully photographed to ensure that the measurement paper was centered in the frame.
Immediately after the images were captured, a preliminary analysis was conducted to verify the completeness of the image processing pipeline. All the measurement papers were checked to ensure that they could be included in the dataset. In cases where the analysis program encountered exceptions, rephotography was performed to prevent bias due to missing information from the measurement paper during analysis. Consequently, this study focused exclusively on analyzing trichome density from images where AR markers could be detected.
The accuracy of AR marker detection was assumed to impact the results inconsiderably. This assumption was based on the practical consideration that the detection accuracy of AR markers and other 2D codes depend primarily on the library used and has already reached a practical level for applications. Moreover, in the event of an error, the kit user can rectify the problem by performing rephotography.
After the data collection and preprocessing steps, information from all the measurement papers was included in the final dataset, which comprised 6,855 images suitable for analysis. This comprehensive dataset encompasses a wide range of distances, angles, and lighting conditions and ensures that the developed model is trained and evaluated on a diverse set of scenarios, closely mimicking real-world usage of the kit.
@@ 3.3.1 Configuration of the data analysis environment
The analytic environment was built using a Linux-based virtualization technology to ensure consistency and reproducibility of the analytical results across different computer systems, regardless of the specific hardware specifications. The analysis was conducted using Python [pythonsoftwarefoundationpython] version 3.10. Table 3 provides a comprehensive list of the Python library versions and operating system versions utilized in this study. The opencv-contrib-python, imbalanced-learn [scikitlearncontribimbalancedlearn], lightgbm [kguolinandmicrosoftresearchlightgbm], shap [slundbergandsleeshap], and ExifRead [sianareexifread] libraries were utilized for image processing, dataset preprocessing, machine learning model training, model interpretation, and metadata extraction from image data, respectively.
@@ 3.3.2 Statistical analysis of nitrate ion concentrations
We employed a series of nonparametric tests to statistically validate the observations of nitrate ion concentrations across different fertilizer levels. First, we used the Kruskal-Wallis test to examine the overall differences in distribution among the three fertilizer levels. Next, we conducted pairwise comparisons via Mann-Whitney U tests to assess the differences between each pair of fertilizer levels. To account for multiple comparisons, we applied the Bonferroni correction to the p values obtained from the Mann-Whitney U tests.
@@ 3.3.3 Multicollinearity and variable interaction assessment
Before building the model, we investigated the potential presence of multicollinearity and interaction effects among the predictor variables. Multicollinearity occurs when predictor variables are strongly intercorrelated, which can result in unstable and unreliable regression coefficient estimates. To detect the presence of multicollinearity, we utilized the variance inflation factor (VIF), which measures the extent to which the variance of a regression coefficient estimate increases owing to the correlation among predictor variables. A VIF value surpassing the cutoff point of 10 is commonly considered a strong indication of problematic multicollinearity [zuur2010a]. To enhance the model interpretability and alleviate the consequences of multicollinearity, we removed variables with VIF values surpassing the threshold of 10 from further analyses. Moreover, we explored the possibility of interaction effects among the predictor variables, which arises when the level of a predictor variable influences the impact of another predictor variable on the response variable. To uncover potential interactions between features, we conducted a stratified analysis involving dividing the data into subgroups on the basis of the levels of a suspected interacting variable and assessing relationships between the predictor and response variables within each subgroup. Specifically, we investigated how image resolution might modulate the association between trichome density and nitrate concentration in tomato leaves.
@@ 3.4.1 Problem formulation and objectives
The primary objective of this study was to develop a kit that is suitable for practical use in the field. By leveraging existing knowledge on the correlation between leaf nitrogen-related indicators, we aimed to address a crucial question for farmers, which is affected by the fertilizer stress level and trichome density: Do my tomato plants in the field require additional fertilization?
@@ 3.4.2 Model design and variable selection
Consequently, we formulated a binary classification problem to determine the need for fertilization on the basis of leaf nitrate ion concentrations. The threshold for the leaf nitrate ion concentration was set as a parameter. Importantly, there is no universally ideal level of fertilizer stress for tomatoes. This level can vary depending on the environmental conditions and specific goals of the cultivator. Recognizing this variability, our study did not set a specific nitrate ion threshold. Instead, we parameterized the threshold for the leaf nitrate ion concentration in our model. This threshold was used to determine whether additional fertilization was necessary: If the concentration exceeded the threshold, no additional fertilization was needed, whereas concentrations below the threshold indicated a need for fertilization. This binary decision was then used as the target variable in our model.
For the explanatory variables in this model, we used the NND, image resolution and exposure time. We applied standard scaling to these explanatory variables to ensure that all variables were on a comparable scale and to improve the model’s performance. This scaling technique transforms the variables so that their mean is zero and their standard deviation is one.
In our cross-validation process, the data were divided into training and test sets. Importantly, the primary focus of this research was to develop a kit suitable for practical field use, with an emphasis on architectural-level improvements. Specifically, a quantitative method for characterizing tomatoes and refinement of measurement methods were selected. Therefore, we did not employ hyperparameter tuning using a validation dataset. Instead, we used default values for the model’s hyperparameters. This decision may lead to underfitting or overfitting due to insufficient control over model complexity, potentially resulting in reduced generalization performance and missed opportunities for theoretical performance analysis and optimization [mohri2018foundations]. However, we deemed these potential drawbacks acceptable, given that our primary objective was to prioritize the development of a practical and accessible kit for field use, rather than achieving the highest possible model performance.
@@ 3.4.3 Cross-validation strategy
To evaluate the model’s generalization performance, we employed leave-one-out cross-validation (LOOCV) for each compound leaf. This method was chosen because of the relatively small-scale dataset, derived from 40 compound leaves harvested from 20 tomato plants from the three intervention groups. These compound leaves yielded 300 leaflets, and we analyzed 6,855 images taken of these leaflets. Given this dataset structure, LOOCV was deemed suitable for minimizing bias from specific samples while maximizing the use of our limited but diverse data points. Furthermore, our feature extraction process enables learning with tabular data, maintaining low computational costs despite the iterative nature of LOOCV.
In each iteration of the LOOCV process, a single unique compound leaf was selected. The data derived from this selected compound leaf served as the test set, whereas the data from the remaining compound leaves constituted the training set.
@@ 3.4.4 Evaluation methodology
Because the kit users can easily and cost-effectively perform multiple measurements, we evaluated the model’s predictive performance via a method that averages multiple predictions obtained from multiple images into a single prediction. We randomly extracted 1 to 25 rows of data from the test set and input them into the model. This method simulates scenarios where users capture and utilize 1 to 25 images simultaneously for prediction, which allows us to assess the model’s performance under various measurement intensities. This assessment enables us to understand how the model’s performance improves with an increasing number of measurements. To ensure a comprehensive evaluation, we aggregated the predictions and actual observations from all cross-validation iterations: This combined data is used to calculate various performance metrics.

### v3

@@ 3.2.2 Tomato growth conditions in hydroponic culture
The nutrient solution in each bucket was continuously aerated and mixed via a blower system (YASUNAGA, AP-40P), which distributed air through eight segments of silicone tubing and dispersed it into the nutrient solution via air stones for proper oxygenation. To prevent algae growth in the nutrient solution, the buckets were fitted with lids featuring center holes for tomato stem placement.
@@ 3.2.3 Fertilizer concentration settings
To investigate the relationship between fertilizer stress and tomato phenotypes, we prepared nutrient solutions using hydroponic fertilizer (OAT Agrio, Fertilizer Series SA) at three concentrations of total dissolved solids (TDS) to induce various levels of fertilizer stress: 1500 ppm (high), 1000 ppm (medium), and 500 ppm (low). This interventional research design enabled us to examine both the presence of intervention effects on tomato phenotypes and their dose-response relationships.
@@
We established the medium and high concentrations at two and three times this baseline. Under our greenhouse conditions, preliminary cultivation experiments confirmed the suitability of this range, showing no visually identifiable growth defects or plant death at 500 ppm and no apparent growth inhibition due to increased osmotic pressure even at 1500 ppm. These results validated our concentration range selection while supporting our primary goal of detecting fertilizer stress through plant phenotypes to provide actionable fertilization recommendations.
@@ 3.2.4 Experimental design and management
To evaluate the effects of fertilizer levels on key outcome variables, i.e., nitrate ion concentration in tomato leaves, trichome density, and yield, we implemented a randomized block design that minimized the potential impact of spatial variations in environmental factors [fisher1966design]. The fertilizer treatment plots were arranged as shown in Figure 7, with the preliminary experiment utilizing the vacant areas depicted in the figure.
Nutrient solution management followed a systematic protocol throughout the cultivation period. Water was added to the buckets when nutrient solution levels decreased, while fertilizer was applied biweekly, with concentrations adjusted to the desired TDS levels via a TDS meter (EUTECH, PCSTestr35). Although the fertilizer composition closely matched the nutrient uptake ratios of tomatoes, enabling operation as a closed hydroponic system without frequent solution replacement [ishihara2007effect], practical considerations necessitated occasional solution replacement when the solution became cloudy due to algae or root debris.
Plant survival was monitored throughout the experiment, with some attrition occurring at different stages. Two plants were lost before the compound leaf harvesting stage, reducing the study population to 20 plants. By the end of the fruit harvesting period, the loss of two additional plants further reduced the population to 18 plants for yield assessment.
@@ 3.2.5 Compound leaf sampling and trichome transfer
Our experimental protocol included systematic collection of compound leaves from tomato plants, with a total of 40 samples collected across two harvesting sessions in March and April 2023. To ensure consistency and minimize potential biases, we implemented a structured sampling process where samples were processed sequentially according to ascending plant identification numbers. These IDs had been randomly assigned to different fertilizer treatment levels during the randomized block design setup, preventing any treatment-related bias in the sampling order.
The sampling location was standardized at 15 cm below the apical meristem of each plant. When a suitable compound leaf was unavailable at this position, we temporarily skipped the plant to allow for additional growth, continuing the sampling process iteratively until all the plants were sampled. For each compound leaf, we selected all leaflets larger than the measurement paper opening for trichome density measurements. The number of measurement papers matched the number of selected leaflets per compound leaf, allowing for accurate assessment of trichome density distribution across each leaflet.
The glandular trichome transfer process followed a simple yet precise protocol that balanced measurement reliability with practical field applicability. Each leaflet was carefully positioned on the adhesive side of the cellophane tape (Figure 8) and gently pressed with a finger to achieve complete contact between the leaf surface and the tape. Following leaflet removal, we photographed the cellophane tape containing the transferred trichomes alongside the measurement paper via a smartphone camera.
@@ 3.2.6 Image acquisition and dataset preparation
Image acquisition was performed via a smartphone (OPPO, Reno A) equipped with a 16-megapixel camera featuring an F1.7 aperture and a CMOS sensor (Sony, IMX398). We utilized the device’s native camera application with its autofocus capability and automatic adjustment of exposure, ISO sensitivity, and focus parameters, leveraging built-in optimization algorithms to ensure clear images across varying conditions.
To ensure a comprehensive dataset that accounts for various usage scenarios, we manually varied the image capture parameters to validate system robustness. The distance between the camera sensor and measurement paper ranged from 79 to 218 mm, with horizontal offset angles ranging from -12.52 to 10.62 degrees and vertical offset angles ranging from -9.04 to 8.13 degrees, simulating natural variations in user handling. The measurement paper was positioned with sufficient background separation to ensure that the background was not clearly identifiable in the captured images. Photography was conducted in a room with a north-facing window, which eliminated direct sunlight and provided diffuse natural light, supplemented by the smartphone’s built-in white LED and oblique ceiling fluorescent lighting (Figure 4). Photographic sessions were scheduled across different times of day, with solar elevation angles ranging from -7.95 to 33.42 degrees.
Quality control measures were implemented throughout the image acquisition process. Each image was composed with the measurement paper centered in the frame, and preliminary analysis was conducted immediately after capture to verify pipeline compatibility. All the measurement papers were carefully checked for dataset inclusion, with immediate rephotography performed when the analysis program encountered exceptions, mostly due to failures in two-dimensional code detection.
The detection accuracy of these codes primarily depends on the code type and detection libraries used. As two-dimensional code technology has already achieved practical reliability levels, users in agricultural field conditions can simply retake photographs when the measurement software indicates detection exceptions. Given this independence between code detection and our primary objectives of trichome density measurement and fertilization recommendations, we excluded code detection accuracy from our verification targets.
After the data collection and preprocessing steps, information from all the measurement papers was compiled into the final dataset, which comprised 6,855 images that successfully passed through the image processing pipeline. This comprehensive dataset encompasses a wide range of distances, angles, and lighting conditions, ensuring that the developed model is trained on and evaluated based on diverse scenarios that closely mimic real-world usage of the kit.
@@ 3.2.7 Nitrate concentration measurement in tomato leaves
Following trichome density measurements, we developed a systematic process for extracting and analyzing leaf nitrate concentrations. The tomato leaves were first frozen via a refrigerant (HFO-1234ze) to disrupt cell walls through freezing, facilitating subsequent juice extraction. The leaves were placed in plastic bags with their zippers partially closed, leaving a small opening for refrigerant introduction. During the freezing process, the evaporating refrigerant generated dry gas that displaced the moist air from the bags, thus preventing frost formation.
[…]
@@ 3.3.1 Configuration of the data analysis environment
The analytic environment was built via container virtualization technology to ensure the consistency and reproducibility of the analytical results across different computer systems, regardless of the specific hardware specifications. The analysis was conducted via Python [pythonsoftwarefoundationpython] version 3.10; Table 3 provides a comprehensive list of the Python library versions and operating system versions utilized in this study. We used opencv-contrib-python for image processing, imbalanced-learn [scikitlearncontribimbalancedlearn] for dataset preprocessing, lightgbm [kguolinandmicrosoftresearchlightgbm] for machine learning model training, shap [slundbergandsleeshap] for model interpretation, and ExifRead [sianareexifread] for metadata extraction from image data.
@@ 3.3.2 Statistical analysis of nitrate ion concentrations
To statistically validate the observations of nitrate ion concentrations across different fertilizer levels, we employed a series of nonparametric tests. First, the Kruskal-Wallis test was used to examine the overall distribution differences among the three fertilizer levels. Next, pairwise Mann-Whitney U tests were conducted to assess the differences between each pair of fertilizer levels, with Bonferroni correction applied to the resulting p values to account for multiple comparisons.
@@ 3.3.3 Multicollinearity and variable interaction assessment
Before model construction, we conducted a comprehensive analysis of the relationships among the predictor variables, specifically examining multicollinearity and interaction effects. Multicollinearity, which occurs when predictor variables are strongly intercorrelated, can lead to unstable and unreliable regression coefficient estimates. We assessed multicollinearity via the variance inflation factor (VIF), which measures the extent to which the variance of a regression coefficient estimate increases owing to the correlation among predictor variables. Following common practice, variables with VIF values exceeding 10 were considered to indicate problematic multicollinearity [zuur2010a], and they were removed from subsequent analyses to enhance model interpretability.
We also examined the interaction effects among the predictor variables, which arise when the level of a predictor variable influences the impact of another predictor variable on the response variable. Our investigation focused particularly on how image resolution might modulate the relationship between trichome density and leaf nitrate concentration. This analysis was conducted through stratification, dividing the data into subgroups on the basis of levels of suspected interacting variables and examining the predictor-response relationships within each subgroup.
@@ 3.4.1 Problem formulation and objectives
The primary objective of this study was to develop a kit suitable for practical use in the field. By leveraging existing knowledge about the correlation between trichome density and leaf nitrogen-related indicators, we aimed to address a crucial question that farmers face, which is affected by the fertilizer stress level and trichome density: Do my tomato plants in the field require additional fertilization?
@@ 3.4.2 Model design and variable selection
We formulated a binary classification problem to determine the need for fertilization on the basis of leaf nitrate ion concentrations. Given that there is no universally ideal level of fertilizer stress for tomatoes, as optimal levels vary with environmental conditions and cultivator goals, we designed our model with a parameterized nitrate ion concentration threshold. Using this threshold, our model makes binary decisions. Concentrations above the threshold indicate that no additional fertilization is needed, whereas concentrations below the threshold suggest that fertilization is needed.
The model used three explanatory variables: NND, image resolution, and exposure time. These variables underwent standard scaling to ensure comparable scales and to improve model performance. This scaling technique transformed each variable to have a mean of zero and a standard deviation of one. For model evaluation, we employed a cross-validation process that divided the data into training and test sets.
The primary focus of this research was to develop a kit suitable for practical field use, emphasizing improvements at the architecture level, that is, the system’s fundamental structure and design. We aimed to develop objective methods for tomato characterization rather than optimizing model performance. Therefore, we used the model’s default hyperparameter values instead of implementing extensive tuning with a validation dataset. While this decision might lead to underfitting or overfitting due to insufficient control over model complexity, potentially reducing generalization performance and limiting the opportunities for theoretical performance analysis [mohri2018foundations], we considered these trade-offs to be acceptable, as our priority was creating an accessible and practical tool rather than maximizing theoretical model performance.
@@ 3.4.3 Cross-validation strategy
Our model validation approach was determined by the hierarchical structure of our dataset. We analyzed 40 compound leaves harvested from 20 tomato plants across three intervention groups, which yielded 300 leaflets and, ultimately, 6,855 analyzed images. Given this relatively small-scale but diverse dataset, we employed leave-one-out cross-validation (LOOCV) at the compound leaf level to evaluate the model’s generalization performance. This approach was selected to minimize bias from specific samples while maximizing the use of our limited data points.
LOOCV implementation was supported by our feature extraction process, which enables learning with tabular data while maintaining low computational costs despite the iterative nature of LOOCV. In each iteration, we selected one compound leaf as the test set and used the data from all remaining compound leaves for training.
@@ 3.4.4 Evaluation methodology
Our evaluation strategy was designed to reflect the practical usage of our kit, where users can easily perform multiple measurements cost-effectively. We developed a method to assess model performance by averaging the predictions from multiple images, simulating realistic usage scenarios where farmers might take between 1 and 25 images for a single assessment. This approach was implemented by randomly extracting 1 to 25 rows of data from the test set and inputting them into the model for prediction. By examining the model’s performance across different numbers of measurements, we could evaluate how prediction accuracy improves with increased measurement intensity.
For a comprehensive performance assessment, we aggregated predictions and actual observations across all cross-validation iterations to calculate various performance metrics. This aggregated evaluation approach provides insights into the model’s reliability under different measurement scenarios while maintaining practical relevance to field applications.